OpenStudy (chosenmatt):

Which values of x and y are solutions to the open sentence? 3y ≥ 2x + 5 A. x = 1, y = 1 B. x = 1, y = 2 C. x = 2, y = 2 D. x = 2, y = 3

OpenStudy (anonymous):

D. just plug it in

OpenStudy (anonymous):

Substitute* them in, sorry.

OpenStudy (chosenmatt):

I still don't get it...

OpenStudy (aum):

\(3y \ge 2x + 5\) Let us try the first choice: A. x = 1, y = 1 put x = 1 and y = 1 in the given inequality: \(3*1 \ge 2*1 + 5\) \(3 \ge 7\) This is NOT true. Try the other three choices.

OpenStudy (anonymous):

Lets try this...

OpenStudy (anonymous):

Ill let aum handle this.

OpenStudy (anonymous):

aum he had trouble understanding the greater thn equal to symbols, I remember form his other problems